QUESTION 30

  1. Consider this question:

    For a certain set of data, what percentage of individuals would have a z-score greater than 1.08?

    (Hint: this is asking what percentage of the normal curve would fall above that z-score.)

    (Hint: think about where this z-score is located in relation to the mean. Think about the direction being asked about - (less than or greater than) that value).

    Which choice below is true?

    A.

    The answer will be 100%.

    B.

    The answer will be 10.8%

    C.

    The answer will be more than 50%

    D.

    The answer will be less than 50%.

    QUESTION 31

  2. Consider this question:

    For a certain set of data, what percentage of individuals would have a z-score greater than -2.20?

    (Hint: this is asking what percentage of the normal curve would fall above that z-score.)

    (Hint: think about where this z-score is located in relation to the mean. Think about the direction being asked about - (less than or greater than) that value).

    Which choice below is true?

    A.

    The answer will be close to 50%.

    B.

    The answer will be about 22%.

    C.

    The answer will be higher than 95%.

    D.

    The answer will be less than 5%.

    QUESTION 32

  3. Consider this question:

    For a certain set of data, what percentage of individuals would have a z-score greater than 2.20?

    (Hint: this is asking what percentage of the normal curve would fall above that z-score.)

    (Hint: think about where this z-score is located in relation to the mean. Think about the direction being asked about - (less than or greater than) that value).

    Which choice below is true?

    A.

    The answer will be higher than 95%.

    B.

    The answer will be about 22%.

    C.

    The answer will be close to 50%.

    D.

    The answer will be less than 5%.
Answer #1

Use Excel function =NORMSDIST (z)

30. P(z> 1.08)= P(z< -1.08)

=NORMSDIST (-1.08)

= 0.1401

= 14.01%

D.

The answer will be less than 50%.

_______________

31.

P(Z>-2.20)= P( z< 2.20) = NORMSDIST (2.20) = 0.9861

so, (C)

________&&&&_&

32.

P(z>2.20)= P(z < -2.20) = 0.0139

So, (D)